### v3

# Sparsity Forcing: Reinforcing Token Sparsity of MLLMs

Abstract
Sparse attention mechanisms aim to reduce computational overhead with minimal accuracy loss by selectively processing salient tokens. Despite their effectiveness, most methods merely exploit a model’s inherent sparsity and thus plateau at moderate budgets (about 50% token reduction), with little headroom to push budget lower without hurting accuracy. Other approaches attempt to enforce sparsity through trainable sparse attention or sharpness-inducing regularizers, but these either fix rigid patterns that ignore input and layer dynamics, or optimize proxy objectives without direct control over token budgets. In this paper, we explicitly reinforce token sparsity in well-posed multimodal large language models (MLLMs) through a simple RL-based post-training framework named Sparsity Forcing. Our method explores the efficiency-accuracy trade-off by running multiple rollouts with different token budgets, where both efficiency (token reduction ratio) and performance (answer correctness) are formulated as joint rewards. By contrasting rollouts within each group, the more efficient and correct answer is rewarded while less efficient or incorrect ones are penalized, thereby turning token saving into an end-to-end, inference-consistent optimization objective. Across thirteen image and video benchmarks, Sparsity Forcing raises token reduction ratio on Qwen2-VL/Qwen2.5-VL from 20% to 75% with minimal accuracy decline, significantly reducing long-context inference memory by up to 3 while speeding up decoding by up to 3.3.
1 Introduction
Multimodal Large Language Models (MLLMs) (Chen et al., 2024c; Li et al., 2024a; Yang et al., 2024a) have achieved impressive results in tasks such as image captioning and visual question answering (Fu et al., 2023; Li et al., 2024b). However, when processing high-resolution images or long videos, the visual encoder often produces an excessive number of visual tokens, severely constraining the generative efficiency of MLLMs (Chen et al., 2024a). Existing sparse attention methods (Chen et al., 2024a; He et al., 2024; Jiang et al., 2024) exploit the sparsity of attention maps to prune redundant tokens, thereby reducing memory and latency. The core idea is that many tokens receive negligible attention and can be safely discarded, as in FastV (Chen et al., 2024a), which drops half of the visual tokens after the second layer with little accuracy loss. However, further budget reduction—for instance, to 20% or even 10% of the total tokens—remains challenging, since these approaches merely leverage naturally emergent sparsity of MLLMs rather than actively enforcing it, which fundamentally affects the achievable inference-time acceleration.
Early attempts to enhance token sparsity typically rely on either new attention architectures or sharpness-inducing regularizers on attention maps. For instance, MOBA (Lu et al., 2025) and NSA (Yuan et al., 2025) enforce semi-structured sparsity through trainable sparse attention in LLMs. However, such approaches predefine rigid sparsity patterns and ignore the dynamics across inputs, layers, and training stages. They also require training from scratch, making them less practical when adapted to MLLMs in a post-training setting. Alternatively, works such as (Gong et al., 2024; Peruzzo et al., 2024; Araabi et al., 2024) introduce structured regularization losses that concentrate attention on semantically relevant regions, thereby reducing token redundancy. Yet these methods optimize proxy objectives like attention sharpness, which does not necessarily translate into consistent end-to-end token savings. Crucially, both streams of work are implemented in the SFT regime under teacher forcing, where sparsity is enforced on ground-truth tokens rather than generated outputs. This creates a mismatch with inference and often limits the realized end-to-end efficiency gains.
In this paper, we propose a reinforcement learning-based post-training method, Sparsity Forcing , that directly optimizes the performance-efficiency trade-off of MLLMs via the Group Relative Policy Optimization (GRPO) (Shao et al., 2024). For each vision-language query, we optimize a joint reward that increases with final-answer correctness and with token reduction ratio, turning token saving into an end-to-end objective rather than a proxy. Specifically, as shown in Figure 1, we treat an MLLM with sparse attention (e.g., Qwen2-VL+ZipVL (Wang et al., 2024; He et al., 2024)) as the policy model and the original MLLM with standard causal attention as the reference model; this anchoring stabilizes learning and preserves task fidelity, limiting accuracy loss even at high sparsity. To discover the smallest budget that preserves correctness, we execute multiple rollouts under adaptive token budgets to generate answers—a progressive budget sweep that dynamically tests whether low-salience tokens are actually needed for answer correctness. For each rollout, we compute the joint efficiency-performance reward, then contrast rewards within the group to define the advantages: rollouts that are both correct and more efficient tend to receive positive advantages, while less efficient or incorrect rollouts receive lower or negative advantages; then the policy is updated accordingly. Over training, this update progressively concentrates computation on the most informative tokens. Compared with previous SFT-based methods, our RL-based method applies an inference-aligned token pruning policy and KV-cache management for deployment, thereby yielding real end-to-end savings on MLLMs with minimal performance decline.
Our contributions can be summarized as follows:
(1) We propose a novel post-training framework named Sparsity Forcing that explicitly promotes token sparsity in MLLMs, aiming to enhance their achievable efficiency during inference.
(2) We cast efficiency-performance as an explicit joint reward rather than a proxy, producing deployment-aligned sparsity while requiring neither architecture changes nor training from scratch.
(3) Experimental results demonstrate that Sparsity Forcing can enhance the sparsity of Qwen2/2.5-VL from 20% to 75% with a minimal performance decline on thirteen benchmarks, significantly reducing long-context inference memory by up to 3 while speeding up decoding by up to 3.3.
2 Related Work
Sparse attention of MLLMs. Sparse attention is designed to alleviate the computational constraints during deployment, which arise from the extended visual sequence. A line of work achieves it through token selection. ZipVL (He et al., 2024) proposed a dynamic ratio allocation strategy of important tokens, which is adaptively determined based on the layer-specific distribution of attention scores. VisionZip (Yang et al., 2024b) is a text-agnostic visual token compressor that selects dominant tokens by attention of the visual encoder and merges the rest by semantic similarity. SparseVLM (Zhang et al., 2024b) selects relevant text “raters” via attention, adaptively prunes visual tokens with a rank-based per-layer ratio, and recycles pruned tokens. SeerAttention (Gao et al., 2024) developed a self-distillation training scheme to efficiently train the AttnGate as a binary classification of token selection. Another line assigns sparsity patterns to attention heads. For example, Minference (Jiang et al., 2024) and MMinference (Li et al., 2025) determined the optimal pattern for each attention head and dynamically built sparse indices based on the assigned pattern during inference. However, these methods primarily leverage inherent attention sparsity at inference rather than actively encouraging token sparsity during optimization, which can lead to accuracy degradation under extremely low budgets. In this paper, we propose Sparsity Forcing, a post-training RL-based framework that explicitly encourages token sparsity by formulating token selection and answer correctness as an efficiency-performance reward-driven decision process that aligns with the actual inference pipeline.
Encouraging token sparsity. Recently, trainable sparse attention methods exploit the sparsity patterns of attention during training, which can also be deemed as encouraging token sparsity toward a predefined budget with end-to-end training. For example, NSA (Yuan et al., 2025) designed a dynamic hierarchical sparse strategy, combining coarse-grained token compression with fine-grained token selection to preserve both global context awareness and local precision. MOBA (Lu et al., 2025) proposed a block-wise attention probing method and reformulated the sparse attention as Mixture of Experts. However, these methods typically overlook the dynamics across inputs, layers, and training stages. They also require training from scratch, which makes them less practical and often suboptimal when directly adapted to well-posed MLLMs in a post-training setting. Another line of work explicitly sharpens attention by adding entropy-minimizing or norm-based regularizers to the attention map, e.g., minimum-entropy/ penalties for NMT (Zhang et al., 2018) and spatial-entropy regularization for ViTs (Peruzzo et al., 2024), which yield sharper, more selective attention distributions. Nevertheless, such sharpening losses optimize a proxy objective: they expose no explicit control over the retained-token budget, and the induced sharpness does not reliably translate into end-to-end token savings or stable accuracy when post-training MLLMs. In this paper, we reformulate the objective as a joint efficiency-performance reward with GRPO and train with multi-budget exploration.
3 Our Method
3.1 Preliminaries
Token-level sparse attention. Let be query, key, and value matrices in the attention mechanism, respectively, where is the sequence length and denotes the embedding dimension. In token-level sparse attention, the attention computation is restricted exclusively to a subset of selected tokens in the query and key matrices. Formally, two binary indicator vectors are introduced to specify the selection, where the value of indicates a selected token, and otherwise. For the query and key matrices, the corresponding masks are constructed as:
where is a vector of ones. The masks then have rows of ones for selected tokens, and zeros elsewhere. Finally, the output of the sparse attention mechanism is computed by:
where is the row-wise Softmax function, and denotes the element-wise matrix multiplication.
The token budget of sparse attention is given by , where is the zero-norm function that counts the number of nonzero elements. In this paper, we adopt top- sparse attention, a mechanism derived from top- (nucleus) sampling, where the smallest set of tokens whose cumulative attention scores exceeds a threshold is retained for subsequent computation. This design enables dynamic adjustment of sparsity across different layers or tasks while providing a theoretical upper bound on the approximation error of (Lin et al., 2025). Consequently, it supports an online efficiency policy during training with controllable resulting error. Mathematically, the query and key masks are tailored to solve the following constrained optimization problem:
where the sum of attention scores equals to due to the row-wise Softmax function.
Inference pipeline with sparse attention. Given a vision-language question , where is the joint input space of multimodal data, MLLMs first prefill the input with sparse attention and save the pruned KV cache into memory for decoding. Ideally, the final answer is expected to be the same as that without using sparse attention. Since the answer is usually short, that is, , the final KV cache usage is approximately . The token ratio is defined as the average proportion of selected tokens per sample, computed as for a sample with token length over all transformer layers .
3.2 Sparsity Forcing
Overview. As shown in Figure 1, we use an MLLM (e.g., Qwen2-VL (Wang et al., 2024)) with sparse attention (ZipVL (He et al., 2024)) as our policy model , and the same model with frozen parameters and standard causal attention as the reference model . For each question , we perform independent rollouts through the policy model to generate a set of distinct responses , where each response is generated with a random threshold , for . In addition, we record the token ratios associated with these responses. Then we compute the reward of each answer with a specially designed reward function that simultaneously considers the performance-efficiency trade-off. As illustrated in Figure 2, sweeping constitutes a progressive test of whether low-salience tokens are necessary: correctness changes across reveal the smallest budget that preserves accuracy. Such a more efficient and correct rollout receive a positive advantage in GRPO, while all other rollouts—either incorrect or less efficient—receive less or even a negative advantage. Finally, we use the corresponding advantages to update the policy model. Here are three key benefits of Sparsity Forcing: (1) We make the efficiency-performance trade-off an end-to-end objective rather than a proxy—GRPO jointly rewards token reduction and answer accuracy. (2) Multi-budget rollouts obtained by sweeping thresholds enable dynamic exploration of the minimum tokens needed for correctness across layers, inputs, and training dynamics, avoiding rigid, hand-crafted positive/negative labels for RL. (3) The training loop mirrors inference: the same token pruning policy and KV-cache management are applied during training and deployment, yielding predictable, budget-aware behavior across tasks and hardware.
Dynamic sparse attention. Our approach can be integrated with any sparse attention mechanism. In this paper, we build up our method on top of ZipVL, which is a typical top- sparse attention. Consider an attention layer with input tokens, where the standard attention score matrix is denoted as . The accumulated attention score for each token is calculated by summing the corresponding column:
These accumulated attention scores are subsequently sorted in descending order, such that represents the -th highest attention score. Then, the number of important tokens is determined by preserving the majority of attention scores with minimal number of tokens, which can be expressed as:
After determining the number of important tokens for each layer, we use normalized attention scores to assess token importance, calculated as follows:
Here, denotes the number of non-zero elements in the -th column. Important tokens are then selected using the top- indexing method, while the remainder are considered less important:
Inference is then carried out using only the important tokens in for each layer, which participate in subsequent self-attention during both prefilling and decoding, while unimportant tokens are skipped.
Reward function. The reward of our method is composed of performance and efficiency rewards. Specifically, the performance reward function assigns a binary reward where a correct answer is assigned to 1 and otherwise 0 (Guo et al., 2025). The efficiency reward is given by the token reduction ratio , where a higher value indicates more extensive pruning of redundant tokens, thus leading to better efficiency. However, a naive additive combination of them becomes efficiency-biased when no rollout in the group is correct, because the accuracy signal collapses to a constant while the efficiency term still pushes budgets downward. To probe the efficiency-accuracy landscape without collapsing to trivial ultra-sparse policies, we promote efficiency only when the group contains at least one correct answer. Concretely, given a sampling group , we define a group-level indicator:
The per-sample reward is then defined as:
Let be the reward of the generated answers, then the advantage of each answer is obtained by normalizing its reward as:
Eventually, the policy update follows the clipped surrogate objective of GRPO (Shao et al., 2024):
where denotes all learnable parameters to be updated; is the discrete uniform distribution from to ; and are the probabilities under the updated and previous policies, respectively; is the clip operator that constrains the input within a narrow range with a small-valued parameter to prevent large policy updates; measures the KL divergence between the updated and reference policies; and is the weight coefficient of the KL divergence term. The first term facilitates better responses while ensuring stable updates through clipping. In addition, the second term penalizes deviation from a reference model, hence minimizing the transfer entropy caused by the sparse attention.
4 Experimental Setup
| Model | Method | Ratio | MME | MMBench | MMStar | ChartQA | TextVQA | OCRBench | MMMU-Pro | Avg. |
| Qwen2-VL-7B | Full | 100 | 2305 | 78.9 | 57.9 | 81.3 | 82.0 | 807 | 32.9 | 70.9 |
| ZipVL | 79.7 | 2294 | 77.5 | 56.0 | 80.3 | 81.4 | 802 | 32.1 | 69.9 | |
| Sparsity Forcing | 23.6 | 2308 | 78.4 | 58.2 | 81.2 | 82.1 | 807 | 32.9 | 70.8 | |
| Qwen2.5-VL-3B | Full | 100 | 2157 | 78.8 | 55.5 | 83.4 | 78.7 | 784 | 31.5 | 69.1 |
| FastV | 52.1 | 1922 | 75.3 | 52.1 | 78.5 | 76.9 | 737 | 30.2 | 65.1 | |
| VisionZip | 50 | 2026 | 76.6 | 53.5 | 80.5 | 76.4 | 721 | 29.2 | 65.8 | |
| ZipVL | 74.1 | 2142 | 78.1 | 54.3 | 81.4 | 77.6 | 776 | 30.2 | 68.0 | |
| Sparsity Forcing | 22.9 | 2146 | 78.4 | 55.5 | 83.2 | 78.6 | 774 | 30.9 | 68.7 | |
| Qwen2.5-VL-7B | Full | 100 | 2303 | 83.9 | 62.2 | 84.0 | 82.9 | 845 | 36.7 | 73.8 |
| FastV | 52.1 | 2115 | 81.9 | 61.2 | 80.2 | 79.6 | 760 | 34.5 | 69.9 | |
| VisionZip | 50 | 2209 | 83.3 | 61.8 | 80.6 | 79.9 | 796 | 34.6 | 71.2 | |
| ZipVL | 79.5 | 2290 | 83.9 | 60.4 | 82.0 | 82.6 | 837 | 36.2 | 72.9 | |
| Sparsity Forcing | 24.7 | 2286 | 84.1 | 62.5 | 83.1 | 82.6 | 847 | 36.7 | 73.6 |
| Model | Method | Ratio | VideoMME | MLVU | VideoMMMU | PerceptionTest | EgoSchema | TempCompass | Avg. |
| Qwen2-VL-7B | Full | 100 | 62.9/68.2 | 66.4 | 42.2 | 61.6 | 66.2 | 67.1 | 62.1 |
| ZipVL | 71.2 | 62.1/67.8 | 64.9 | 42.1 | 61.3 | 66.0 | 66.8 | 61.6 | |
| Sparsity Forcing | 23.8 | 61.8/68.3 | 66.0 | 42.4 | 61.3 | 66.2 | 67.3 | 61.9 | |
| Qwen2.5-VL-3B | Full | 100 | 61.2/66.8 | 67.3 | 41.8 | 66.4 | 64.2 | 64.0 | 61.7 |
| ZipVL | 69.4 | 60.6/66.5 | 67.3 | 41.4 | 65.2 | 63.7 | 63.9 | 61.2 | |
| Sparsity Forcing | 22.8 | 61.0/66.5 | 67.2 | 41.9 | 66.2 | 64.0 | 63.7 | 61.5 | |
| Qwen2.5-VL-7B | Full | 100 | 64.5/71.1 | 69.7 | 46.3 | 69.0 | 64.3 | 71.2 | 65.2 |
| ZipVL | 70.3 | 64.1/70.8 | 69.5 | 45.9 | 68.9 | 64.0 | 70.6 | 64.8 | |
| Sparsity Forcing | 25.9 | 64.0/71.0 | 69.5 | 46.2 | 68.7 | 64.2 | 70.9 | 64.9 | |
| LLaVa-Video-7B | Full | 100 | 64.5/70.8 | 70.4 | 35.2 | 66.1 | 57.0 | 66.1 | 61.4 |
| FastV | 51.6 | 64.0/70.5 | 69.7 | 34.8 | 65.8 | 56.2 | 65.4 | 60.9 | |
| VisionZip | 50 | 42.2/60.6 | 52.9 | 31.7 | 42.0 | 39.7 | 46.2 | 45.0 | |
| Minference | 46.1 | 64.3/70.6 | 70.5 | 34.8 | 65.9 | 56.8 | 66.0 | 61.3 | |
| ZipVL | 45.2 | 64.3/70.3 | 70.2 | 34.8 | 65.2 | 56.7 | 65.7 | 61.0 | |
| Sparsity Forcing | 29.6 | 64.1/70.2 | 70.3 | 35.2 | 66.0 | 57.0 | 66.2 | 61.3 |
Implementation. We adopt QwenVL-series models (Qwen2-VL-7B (Wang et al., 2024), Qwen2.5-VL-3B (Yang et al., 2024a), and Qwen2.5-VL-7B) and LLaVA-series models (LLaVA-Video-7B (Zhang et al., 2024c)) as base MLLMs, owing to their strong adaptation to recent R1-based projects. For QwenVL-series calibration, we employ the Video-R1-260k (Feng et al., 2025) dataset, which covers both image and video modalities to ensure broad domain generalization. For LLaVA-Video, we use a 10k subset of LLaVA-Video-178k (Zhang et al., 2024c) to maintain its capability in the video domain. The parameter is set within the range of [0.94, 0.975] with a step size of 0.005 for training. We train our model on 8 NVIDIA A100 GPUs, limiting the number of video frames to 8 for efficiency. The Adam optimizer is used with a learning rate of 1e-6. In the KL divergence term of the GRPO algorithm, the hyperparameter is set to 0.04. During inference, Qwen2.5-VL and Qwen2-VL are evaluated on video benchmarks with approximately 120 frames corresponding to 16,384 tokens, while LLaVA-Video is evaluated with 110 frames corresponding to 20,240 tokens. Notably, during training we employ multiple-budget rollouts to explicitly probe the efficiency-performance trade-off, so that the model learns to enhance token sparsity. At inference, however, we simply fix to 0.975, which is the upper bound of the optimization range, to guarantee accuracy while still reaping the efficiency gains learned during training.
Benchmarks. To assess the effectiveness of our approach, we evaluate on 13 benchmarks: 7 image tasks (MME (Fu et al., 2023), MMBench (Liu et al., 2024a), MMStar (Chen et al., 2024b), ChartQA (Masry et al., 2022), TextVQA (Singh et al., 2019), OCRBench (Liu et al., 2024c), MMMU-Pro (Yue et al., 2024)) and 6 video tasks (VideoMME (Fu et al., 2024), MLVU (Zhou et al., 2024), VideoMMMU (Hu et al., 2025), PerceptionTest (Patraucean et al., 2023), Egoschema (Mangalam et al., 2023), TempCompass (Liu et al., 2024b)). We additionally use HallusionBench (Guan et al., 2024) to assess robustness under low token budgets. Together, these benchmarks span diverse modalities, domains, and difficulty levels for a rigorous evaluation of both efficiency and accuracy. The results are evaluated using LMMs-Eval (Zhang et al., 2024a).
Baselines of enhancing token sparsity. We compare our Sparsity Forcing with the existing two baselines to verify its effectiveness in the post-training scenario. Trainable sparse attention: We adopt MOBA (Lu et al., 2025) and ZipVL into existing MLLMs to explicitly enforce token sparsity toward a predefined budget/attention retention ratio. The block size of MOBA is set to 256, and the budget is set to 25% for salient token selection. Attention sharpness loss: We implement a sharpness-inducing regularization term on attention maps (Gong et al., 2024), encouraging the model to focus its attention distribution on a smaller set of highly relevant tokens while suppressing low-importance ones. We adopt the block-wise attention map probing strategy as MOBA to obtain the attention map. During inference, we select 25% salient tokens for sparse attention.
5 Main Results
Performance comparison with training-free sparse attention. As a training-based enhancement of ZipVL, we first compare against ZipVL. As shown in Tables 1 and 2, our Sparsity Forcing can further reduce the token retention ratio of ZipVL from approximately 80% to 25% in QwenVL-series MLLMs and from 45% to 29% on LLaVA-Video across 13 image and video benchmarks with minimal performance decline. Moreover, compared to other training-free sparse attention methods, Sparsity Forcing achieves a much lower token ratio while maintaining accuracy comparable to the standard-attention baseline, thus demonstrating a superior balance between efficiency and performance. On LLaVA-Video-7B, it matches Minference’s (Jiang et al., 2024) accuracy while using 29.6% tokens vs. 46.1% for Minference. Moreover, it clearly outperforms VisionZip (Yang et al., 2024b) and FastV (Chen et al., 2024a), which both require around 50% tokens but achieve lower or similar accuracy. These results show that Sparsity Forcing can significantly reduce computation by enhancing token sparsity while maintaining performance across image and video benchmarks.
| Method | Ratio | MME | MMStar | ChartQA | VideoMME | Avg |
| Full | 100 | 2303 | 62.2 | 84.0 | 64.5 | 70.2 |
| MOBA | 25.0 | 1906 | 58.6 | 77.3 | 62.6 | 66.2 |
| ZipVL | 61.7 | 2264 | 62.0 | 78.9 | 64.2 | 68.4 |
| Sharpness loss | 25.0 | 1965 | 59.6 | 77.0 | 63.7 | 66.8 |
| Sparsity Forcing | 26.4 | 2286 | 62.5 | 83.1 | 64.0 | 70.0 |
Performance comparison with post-training baseline methods in enhancing token sparsity. As shown in Table 3, Sparsity Forcing achieves a superior efficiency-performance trade-off compared to baseline methods on Qwen2.5-VL-7B. The standard-attention model reaches an average score of 70.0, while MOBA and sharpness loss, both operating at a 25% token ratio, suffer from noticeable drops to 66.2 and 66.8, respectively. ZipVL provides a moderate balance, obtaining 68.4 at a higher token ratio of 61.7%. In contrast, Sparsity Forcing attains 70.3 even with only 26.4% tokens, outperforming all baseline approaches on tasks such as MMStar (Chen et al., 2024b) and ChartQA (Masry et al., 2022). These results highlight the effectiveness of our RL-based optimization that reformulates the efficiency-performance tradeoff as explicit rewards.
Adjustment under different budgets. We further evaluate Sparsity Forcing under varying token budgets to illustrate its adaptability. As shown in Figure 3, Sparsity Forcing enables flexible budget adjustment across tasks and models. On commonsense benchmarks (Liu et al., 2024a), it maintains strong performance even under an extremely low budget of about 10% tokens. On OCR-related tasks (Liu et al., 2024c; Masry et al., 2022), it can further compress the token ratio while still preserving better accuracy than other methods, demonstrating its ability to focus on fine-grained textual regions. Moreover, this property generalizes well across both image and video benchmarks and different model backbones, consistently outperforming other sparse attention baselines under the same token ratio.
Besides, we further discuss the influence of the attention retention ratio during inference. As shown in Figure 4 (a), increasing smoothly raises the token ratio while also improving accuracy, even for the value outside the optimization range, showing that Sparsity Forcing enables controllable efficiency-accuracy trade-offs without abrupt performance drops.
Runtime speed comparison. As shown in Figure 4 (c)(d), Sparsity Forcing delivers clear efficiency gains over existing methods. It achieves up to 3.3 faster inference and 3.0 lower memory usage than FlashAttention-2 (Dao, 2023) at 200k sequence lengths. These results highlight that Sparsity Forcing not only reduces token budgets but also substantially improves runtime efficiency for large-scale multimodal inference.
6 Ablation Study and Discussion
Sparsity Forcing with different sparse attention mechanisms. Table 5 compares Sparsity Forcing using various sparse attention strategies, including top- sampling, threshold-based pruning, and dynamic top- sampling. While top- and threshold-based methods can reduce token usage, they rely on fixed sparsity patterns or constraints and thus act as offline policies, which may not adapt optimally to input-specific variations. In contrast, top- sampling enables online adjustment of token retention according to attention score distributions, achieving the best trade-off between efficiency and performance (e.g., 24.1% token ratio with minimal accuracy drop). This highlights the advantage of coupling grouped-rollout RL with a dynamic, inference-time controllable sparse attention mechanism.
Token sparsity change. Figure 5 shows how training reshapes both the attention distribution and the retained token ratio. After training, the model concentrates attention on a more compact subset of tokens, as illustrated in the left and center figures, demonstrating effective pruning. The layer-wise curves (right) reveal substantial variation in attainable sparsity across depth, indicating that a single global retention rate is suboptimal and a dynamic, layer-aware policy is preferable.
Robustness using a low token budget. Aggressive token pruning can drop essential evidence and increase hallucination risk. To assess robustness, we conduct a sensitivity analysis on HallusionBench (Guan et al., 2024) under a strict token budget. As shown in Figure 6, Sparsity Forcing closely tracks the original model across all accuracy, figure accuracy, and question pair accuracy (aAcc, fAcc, and qAcc) with only minor fluctuations, indicating that our dynamic sparse attention preserves key evidence and does not amplify hallucinations in low-budget settings.
Scalable token sparsity. We examine how learned token sparsity scales with model capacity and context length. As shown in Tables 1 and 2, we find models of different sizes and capabilities in the same MLLM family exhibit broadly similar retained token ratios, a phenomenon that holds across both image and video domains. Moreover, in Figure 4 (b), as the input sequence grows from 4k to 20k tokens, the retained-token ratio steadily decreases while accuracy remains nearly unchanged. This suggests that learned token sparsity adapts to longer sequences by safely discarding more redundant tokens yet preserving essential information, making it particularly effective for high-complexity, long-context video understanding.
| Model | Sparse attention | Ratio | MME | VideoMME |
| Qwen2.5-VL-7B | Full | 100 | 2303 | 64.5 |
| Top- | 25 | 2160 | 60.2 | |
| Threshold | 37.8 | 2218 | 61.6 | |
| Top- | 24.1 | 2286 | 64.0 |
| Range () | Group size | Ratio | MME | VideoMME |
| 8 | 24.1 | 2286 | 64.0 | |
| 12 | 23.7 | 2290 | 63.9 | |
| 8 | 29.0 | 2285 | 63.6 | |
| 12 | 25.7 | 2285 | 63.7 |
Why grouped-rollout RL? Our method requires comparing multiple budgets per query to explore the efficiency-performance tradeoff, where the notion of “positive vs. negative” is inherently non-stationary: as training progresses, sparsity patterns and the minimal correct budget shift, so predefined preference pairs of DPO (Rafailov et al., 2023) and rigid sparse pattern of trainable sparse attention quickly become misaligned with the current trade-off frontier. In contrast, grouped-rollout RL methods (e.g., GRPO, DAPO (Yu et al., 2025)) operate on-policy and define relative advantages within each sampled group of rollouts. This design naturally provides exploration and credit assignment directly tied to budget choices, and avoids maintaining a large, ever-updating preference dataset.
Budget sampling strategy. As shown in Table 5, the ablation of the range of and group size can be interpreted as probing the efficiency-performance trade-off through grouped rollouts. Increasing the group size enables denser coverage of candidate token budgets, allowing the model to explore the budget–accuracy landscape more thoroughly and providing stronger contrastive signals for learning sparsity. However, this benefit comes at the cost of higher training overhead, since more rollouts must be generated per sample. In addition, setting the range relatively high improves the likelihood that at least one rollout within the group is both correct and efficient, which is crucial for contrastive learning to effectively guide sparsity.
7 Conclusion
In this paper, we presented Sparsity Forcing, an RL-based post-training method that explicitly encourages token sparsity in well-posed MLLMs, delivering substantial inference-time acceleration. We treat the token-reduction ratio as the efficiency reward and answer accuracy as the performance reward. Concretely, Sparsity Forcing performs multi-budget rollouts and applies GRPO with group-wise advantages to favor the smallest budget that still yields a correct answer. Compared with previous SFT-based methods, our RL-based approach enforces an inference-aligned token pruning policy and KV-cache management, thereby delivering real end-to-end efficiency gains in MLLMs while keeping performance degradation minimal. Experiments across thirteen image and video benchmarks show that Sparsity Forcing increases the token reduction ratio on Qwen2-VL/Qwen2.5-VL from 20% to as high as 75% with minimal accuracy loss, while cutting long-context inference memory by up to and speeding up decoding by up to .
Limitations and future work. Our current multiple-budget rollout scheme executes answers sequentially for each because every budget induces a different retention mask and KV-cache trajectory, which precludes straightforward batching and slows training. In our future work, we will address this limitation to enable easier and more efficient training. Moreover, beyond token budgets, we also plan to extend reinforced efficiency to broader applications and objectives—such as hardware-aware targets (latency/memory/energy), tool-/retrieval-call budgets, head/layer/MoE expert gating, and KV/activation quantization and retention.